# Discrete equations from Bäcklund transformations of the fifth Painlevé equation

### v1

@@ 4.1 Generalised Laguerre polynomials
Rational solutions in case (i) of Theorem 4.1 are expressed in terms of generalised Laguerre polynomials, which are written in terms of a determinant of Laguerre polynomials, or equivalently as a Wronskian, see [12, §3]. These rational solutions are special cases of the solutions of PV (1.2) expressible in terms of the Kummer functions $M(a,b,z)$ and $U(a,b,z)$ discussed in [43, 57] since
@@
see [54, equation (13.6.19)]. First we define the generalised Laguerre polynomial $T_{m,n}^{(\mu)}(z)$; properties of it are discussed in [6, 12].
@@
See Clarkson and Dunning [12]. The solutions when either $m=0$ or $n=0$ follow from recurrence relations for the Laguerre polynomial, cf. [54, §18.9]. ∎
@@ 4.2 Generalised Umemura polynomials
    The generalised Laguerre polynomial $T_{m,n}^{(\mu)}(z)$ is a Wronskian involving one sequence of Laguerre polynomials, whereas the generalised Umemura polynomial $U_{m,n}^{(\kappa)}(z)$ involves two sequences of Laguerre polynomials, unless either $m=0$ or $n=0$. This is analogous to the situation for rational solutions of PIV (1.10), where one set of rational solutions involves one sequence of Hermite polynomials and a second set involves two sequences of Hermite polynomials, cf. [11, 52].
@@
    When $m=0$, the polynomial $U_{0,n}^{(\kappa)}(z)$ is equivalent to the polynomial defined by Umemura [70] through a differential-difference equation. Also $U_{0,n}^{(\kappa)}(z)$ is equivalent to the determinant $\tau_{m}^{(n)}(z)$ defined in Theorem 5.1; see also [50]. Due to the symmetry (4.19), the same is true for the polynomial $U_{m,0}^{(\kappa)}(z)$.
@@
See Masuda, Ohta and Kajiwara [45, Theorem 1.1].∎
@@ 5.3 Non-unique solutions of the discrete equations
Kitaev, Law and McLeod [40, Theorem 1.2] prove that if $\gamma\not\in\mathbb{Z}$ in case (i), then the rational solution of PV is unique and if $\gamma\in\mathbb{Z}$ then there are at most two rational solutions. Clarkson and Dunning [12] gave examples of non-unique rational solutions which arise for solutions expressed in terms of generalised Laguerre polynomials discussed in §5.1. For example, the rational solutions
@@
5.3.1 First discrete equation: asymmetric dPII
@@ 5.3.2 Second discrete equation
5.3.3 Third discrete equation: ternary dPI
@@ 6 Discussion
Exact solutions of the discrete equations (3.7), (3.17) and (3.27) in terms of Kummer functions and modified Bessel functions (which are a special case of Kummer functions, cf. [54, §13.6(iii)]), using the explicit relationship between these equations and PV (1.2) given here, is currently under investigation, and we do not pursue this further here.
@@ A.1 Properties of the generalised Laguerre polynomials
Clarkson and Dunning [12, §3] prove the following differential-difference equations for the generalised Laguerre polynomials
@@ A.2 Properties of generalised Umemura polynomials
where ${\rm D}_{z}(f{\,\scriptscriptstyle\bullet\,}g)$ is the Hirota operator (4.15), see equation (4.16) and Proposition 4.1 in [45]. From equations (A.12) and (A.13) in [44, Appendix A] one can derive the following differential-difference equations for the generalised Umemura polynomials

### v2

@@ 4.1 Generalised Laguerre polynomials
Rational solutions in case (i) of Theorem 4.1 are expressed in terms of generalised Laguerre polynomials, which are written in terms of a determinant of Laguerre polynomials, or equivalently as a Wronskian, see [16, §3]. These rational solutions are special cases of the solutions of PV (1.2) expressible in terms of the Kummer functions $M(a,b,z)$ and $U(a,b,z)$ discussed in [51, 65] since
@@
see [62, equation (13.6.19)]. First we define the generalised Laguerre polynomial $T_{m,n}^{(\mu)}(z)$; properties of it are discussed in [6, 16].
@@
See Clarkson and Dunning [16]. The solutions when either $m=0$ or $n=0$ follow from recurrence relations for the Laguerre polynomial, cf. [62, §18.9]. ∎
@@ 4.2 Generalised Umemura polynomials
    The generalised Laguerre polynomial $T_{m,n}^{(\mu)}(z)$ is a Wronskian involving one sequence of Laguerre polynomials, whereas the generalised Umemura polynomial $U_{m,n}^{(\kappa)}(z)$ involves two sequences of Laguerre polynomials, unless either $m=0$ or $n=0$. This is analogous to the situation for rational solutions of PIV (1.11), where one family of rational solutions involves one sequence of Hermite polynomials and a second family involves two sequences of Hermite polynomials, cf. [14, 60].
@@
    When $m=0$, the polynomial $U_{0,n}^{(\kappa)}(z)$ is equivalent to the polynomial defined by Umemura [79] through a differential-difference equation. Also $U_{0,n}^{(\kappa)}(z)$ is equivalent to the determinant $\tau_{m}^{(n)}(z)$ defined in Theorem 5.1 below; see also [58]. Due to the symmetry (4.19), the same is true for the polynomial $U_{m,0}^{(\kappa)}(z)$.
@@
See Masuda, Ohta and Kajiwara [53, Theorem 1.1].∎
@@ 5.3 Non-unique solutions of the discrete equations
Kitaev, Law and McLeod [46, Theorem 1.2] prove that if $\gamma\not\in\mathbb{Z}$ in case (i), then the rational solution of PV is unique and if $\gamma\in\mathbb{Z}$ then there are at most two rational solutions. Clarkson and Dunning [16] gave examples of non-unique rational solutions which arise for solutions expressed in terms of generalised Laguerre polynomials discussed in §5.1. For example, the rational solutions
@@
5.3.1 First discrete equation: asymmetric discrete PII
@@ 5.3.2 Second discrete equation
5.3.3 Third discrete equation: ternary discrete PI
@@ 6 Discussion
Tokihiro, Grammaticos and Ramani [78] also derive the discrete system
$x_{n}+x_{n-1}=\frac{1}{y_{n}}+\frac{n+\lambda+\mu}{1-y_{n}},\qquad y_{n+1}y_{n}=\frac{x_{n}-n-\lambda}{x_{n}^{2}-\varphi^{2}},$ (6.1)
with $\lambda$, $\mu$ and $\varphi$ constants, which was derived earlier by Fokas, Grammaticos and Ramani [23] from the Schlesinger transformations of PV (1.2). Tokihiro, Grammaticos and Ramani [78] state that whilst $y_{n}$ satisfies an equation which is equivalent to PV, $x_{n}$ satisfies a second-order, fourth-degree equation. Schlesinger transformations of PV (1.2) are transformations which are quadratic in $\frac{\mathrm{d}{w}}{{\mathrm{d}{z}}}$. For example, consider the Schlesinger transformations $R_{1}$ and $R_{2}$ of PV in [26, 54], which Fokas, Grammaticos and Ramani [23] use to derive the discrete system (6.1). Suppose that $w_{0}=w(z;a_{0},b_{0},c_{0})$ satisfies PV for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}a_{0}^{2},-\tfrac{1}{2}b_{0}^{2},c_{0}\big)$, then the Schlesinger transformations $R_{1}$ and $R_{2}$ are given by
$w_{1}$ $=R_{1}(w_{0})=\frac{\big\{zw_{0}^{\prime}+a_{0}w_{0}^{2}-(a_{0}+b_{0}+z)w_{0}+b_{0}\big\}\big\{zw_{0}^{\prime}+a_{0}w_{0}^{2}-(a_{0}-b_{0}+z)w_{0}-b_{0}\big\}}{z^{2}\big(w_{0}^{\prime}\big)^{\!2}-2a_{0}zw_{0}(w_{0}-1)w_{0}^{\prime}+(a_{0}^{2}w_{0}^{2}-b_{0}^{2})(w_{0}-1)^{2}-2z(a_{0}-c_{0}+1)w_{0}(w_{0}-1)-z^{2}w_{0}^{2}},$
$w_{2}$ $=R_{2}(w_{0})=\frac{\big\{zw_{0}^{\prime}-a_{0}w_{0}^{2}+(a_{0}+b_{0}+z)w_{0}-b_{0}\big\}\big\{zw_{0}^{\prime}-a_{0}w_{0}^{2}+(a_{0}-b_{0}+z)w_{0}+b_{0}\big\}}{z^{2}\big(w_{0}^{\prime}\big)^{\!2}+2a_{0}zw_{0}(w_{0}-1)w_{0}^{\prime}+(a_{0}^{2}w_{0}^{2}-b_{0}^{2})(w_{0}-1)^{2}-2z(a_{0}-c_{0}-1)w_{0}(w_{0}-1)-z^{2}w_{0}^{2}},$
with $^{\prime}=\displaystyle\frac{\mathrm{d}{}}{{\mathrm{d}{z}}}$, and the effect on the parameters is
$R_{1}\big(a_{0},b_{0},c_{0}\big)=\big(a_{0}+1,b_{0},c_{0}-1\big),\qquad R_{2}\big(a_{0},b_{0},c_{0}\big)=\big(a_{0}-1,b_{0},c_{0}+1\big).$
Hence $w_{1}$ and $w_{2}$ respectively satisfy PV (1.2) for the parameters $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(a_{0}+1)^{2},-\tfrac{1}{2}b_{0}^{2},c_{0}-1\big)$ and $(\alpha,\beta,\gamma)=\big(\tfrac{1}{2}(a_{0}-1)^{2},-\tfrac{1}{2}b_{0}^{2},c_{0}+1\big)$, so that $R_{1}R_{2}=R_{2}R_{1}=\mathcal{I}$, the identity transformation. In terms of the Bäcklund transformations $\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}$ defined in §2, it is straightforward to show that
$R_{1}\equiv\mathcal{T}_{1,1,1}^{2}\circ\mathcal{T}_{-1,-1,-1},\qquad R_{2}\equiv\mathcal{T}_{-1,-1,-1}^{2}\circ\mathcal{T}_{1,1,1}.$
All the Schlesinger transformations of PV (1.2) given in [26, 54] can be written as a sequence of the Bäcklund transformations $\mathcal{T}_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}}$. The Schlesinger transformations in [26, 54] are written in terms of the monodromy parameters $\theta_{0}$, $\theta_{1}$ and $\theta_{\infty}$ which are related to the parameters $a_{0}$, $b_{0}$ and $c_{0}$ through (2.5), i.e.
$\theta_{0}=\tfrac{1}{2}(a_{0}+b_{0}-c_{0}-1),\qquad\theta_{1}=-\tfrac{1}{2}(a_{0}+b_{0}+c_{0}-1),\qquad\theta_{\infty}=a_{0}-b_{0},$
so that
$a_{0}=\tfrac{1}{2}(\theta_{0}-\theta_{1}+\theta_{\infty}),\qquad b_{0}=\tfrac{1}{2}(\theta_{0}-\theta_{1}-\theta_{\infty}),\qquad c_{0}=1-\theta_{0}-\theta_{1}.$
Exact solutions of the discrete equations (3.7), (3.18) and (3.28) in terms of Kummer functions and modified Bessel functions, which are a special case of Kummer functions, cf. [62, §13.6(iii)], using the explicit relationship between these discrete equations and PV (1.2), is currently under investigation, and we do not pursue this further here.
@@ Acknowledgements
Conflict of interest statement
The authors declare no conflicts of interest.
Data availability statement
No data were used for the research described in the article.
@@ A.1 Properties of the generalised Laguerre polynomials
Clarkson and Dunning [16, §3] prove the following differential-difference equations for the generalised Laguerre polynomials
@@ A.2 Properties of generalised Umemura polynomials
where ${\rm D}_{z}(f{\,\scriptscriptstyle\bullet\,}g)$ is the Hirota operator (4.15), see equation (4.16) and Proposition 4.1 in [53]. From equations (A.12) and (A.13) in [52, Appendix A] one can derive the following differential-difference equations for the generalised Umemura polynomials